OpenStudy (anonymous):

Using graph paper, graph the following equation. Then click on the graph until the correct one is displayed. 3x - y - 1 = 0

OpenStudy (anonymous):

ive tried but i cant get it

OpenStudy (anonymous):

this is what it would look like

OpenStudy (anonymous):

You need to get it into the form \[y=mx+b\] so just rearrange the equation it gave you 3x-y-1=0 -y-1=-3x -y=-3x+1 y=3x-1

OpenStudy (anonymous):

was my graph right?

OpenStudy (anonymous):

@nocontact115 probably

OpenStudy (anonymous):

@nocontact115 Yes it was

OpenStudy (anonymous):

@Zacharymartin Do you get it now?